Suppose you are given the following Supply and Demand equations. P=80-Q P=20+2Q How much would producers...

  1. Suppose you are given the following Supply and Demand equations.

P=80-Q

P=20+2Q

How much would producers be willing to pay lobbyists to get the government to establish a price floor at $70?
Answer #1

At equilibrium, Qd=Qs

80-Q=20+2Q

Q*= 60/3=20

P*= 80-20=$60

Producer would be willing to pay the lobbyists equap to the rise in producer surplus due to price floor of $70.

Old Producer surplus=0.5(60-20)(20)= 0.5*40*20=$400

When P= $70, Qs= (70-20)/2=25

New producer surplus=0.5(70-20)(25)=0.5*50*25=$625

Rise in producer surplus= Price paid to lobbyists= $625-$400=$225